# Learning Global Hypothesis Space for Enhancing Synergistic Reasoning Chain

### v1

@@ Abstract
Chain-of-Thought (CoT) has been shown to significantly improve the reasoning accuracy of large language models (LLMs) on complex tasks. However, due to the autoregressive, step-by-step generation paradigm, existing CoT methods suffer from two fundamental limitations. First, the reasoning process is highly sensitive to early decisions: once an initial error is introduced, it tends to propagate and amplify through subsequent steps, while the lack of a global coordination and revision mechanism makes such errors difficult to correct, ultimately leading to distorted reasoning chains. Second, current CoT approaches lack structured analysis techniques for filtering redundant reasoning and extracting key reasoning features, resulting in unstable reasoning processes and limited interpretability. To address these issues, we propose GHS-TDA. GHS-TDA first constructs a semantically enriched global hypothesis graph to aggregate, align, and coordinate multiple candidate reasoning paths, thereby providing alternative global correction routes when local reasoning fails. It then applies topological data analysis based on persistent homology to capture stable multi-scale structures, remove redundancy and inconsistencies, and extract a more reliable reasoning skeleton. By jointly leveraging reasoning diversity and topological stability, GHS-TDA achieves self-adaptive convergence, produces high-confidence and interpretable reasoning paths, and consistently outperforms strong baselines in terms of both accuracy and robustness across multiple reasoning benchmarks.
@@ 1 Introduction
LLMs have shown remarkable potential in tasks such as logical reasoning, mathematical proof, and multi-hop question answering. Among them, CoT prompting (Wei et al., 2022) has been demonstrated to improve interpretability and accuracy by decomposing complex problems into coherent intermediate steps. Despite these advances, CoT and its extensions still face critical limitations.
On the one hand, structured approaches such as Tree-of-Thought (ToT) (Yao et al., 2023a), Graph-of-Thought (GoT) (Besta et al., 2024), and Atom-of-Thought (AoT) (Teng et al., 2025) expand the reasoning space beyond single-path CoT and enrich reasoning diversity. However, they lack mechanisms for global integration and interaction across hypotheses, which limits evidence reuse, complicates conflict resolution, and leaves reasoning largely driven by local heuristics. As a result, these methods often fail to achieve systematic semantic integration and remain prone to logical divergence and inconsistency in complex tasks. On the other hand, several systematic frameworks have been proposed to improve efficiency and robustness. For example, ReAct (Yao et al., 2023b) integrates reasoning with acting for interactive tasks, AFlow (Zhang et al., 2024) employs Monte Carlo tree search for reasoning workflows, and ReCEval (Prasad et al., 2023) evaluates reasoning chains for correctness and informativeness. While these approaches provide performance gains, they mostly focus on task-specific outcomes and lack a unified analytical perspective to characterize structural properties of reasoning chains, such as connectivity, cyclicity, and consistency.
Although these methods have achieved certain improvements in performance and efficiency, reasoning chains, as explicit representations of the reasoning process, inherently encode the semantic and logical organization of problem solving and thus hold intrinsic value for structural analysis. However, the absence of a unified analytical framework to systematically characterize the structural properties of reasoning chains makes it difficult to comprehensively evaluate and enhance the reliability, robustness, and interpretability of reasoning outcomes.
@@ 5 Conclusion
Reproducibility Statement
To ensure the reproducibility of our work, we have undertaken systematic efforts in multiple aspects:
Methodological Details. In Section 3, we provide a comprehensive description of the two-phase GHS-TDA framework, including the construction of the Global Hypothesis Graph (GHG) and the subsequent topological analysis (Persistent Homology). Relevant mathematical definitions, formulas, and pseudocode are presented in the main text (Equations (1)–(11)) and Appendix G, enabling precise reproduction of the algorithmic workflow.
Implementation and Parameter Settings. Implementation details are documented in the “Implementation Notes” section and Appendix E. These include the embedding model employed (text-embedding-3-large), the topological analysis toolkit (GUDHI), random seeds (seeds = 5), decoding parameters (temperature = 0.7, top-p = 0.95), the maximum number of path samples (16 LLM calls), as well as default values and tuning strategies for key hyperparameters (e.g., node merging threshold $\theta_{\text{merge}}$, distance weighting factors $\alpha/\beta/\gamma$, number of significant features $K$, and cycle embedding threshold $\delta$).
Datasets and Experimental Configuration. All experiments are conducted on publicly available benchmark datasets (GSM8K, MATH, OlympiadBench, BBH, MMLU-CF, LongBench, HotpotQA, MuSiQue), following their official splits and license agreements. Dataset versions and license information are documented in the appendix to ensure that other researchers can access the same resources.
Verifiability and Interpretability. We provide interpretable representations of reasoning chains (skeleton paths and critical loops), along with cross-dataset visualizations in Appendix D. These materials allow independent researchers to verify whether intermediate reasoning processes are consistent with our reported results.
Supplementary Materials. The supplementary materials include the complete prompts and environment specifications, which facilitate the replication of experiments and further verification of results.

### v2

# arXiv:2602.09794v2[cs.AI] 28 Feb 2026

@@ Abstract
Chain-of-Thought (CoT) has been shown to significantly improve the reasoning accuracy of large language models (LLMs) on complex tasks. However, due to the autoregressive, step-by-step generation paradigm, existing CoT methods suffer from two fundamental limitations. First, the reasoning process is highly susceptible to early-stage errors, which tend to propagate and amplify without a global coordination and correction mechanism, thereby distorting the overall reasoning chain. Second, current CoT methods lack structured analytical frameworks for pruning redundant reasoning and identifying critical reasoning features, resulting in instability and reduced interpretability. To address these issues, we propose Global Hypothesis Structure via Topological Data Analysis (GHS-TDA), which constructs a semantically enriched global hypothesis graph that integrates and coordinates multiple candidate reasoning paths, thereby supporting global consistency refinement and error mitigation. GHS-TDA applies persistent homology-based topological data analysis to capture stable multi-scale structures, remove redundancy and inconsistencies, and extract a more reliable reasoning skeleton. By jointly leveraging reasoning diversity and topological stability, GHS-TDA achieves self-adaptive convergence, produces high-confidence and interpretable reasoning paths, and consistently outperforms strong baselines in terms of both accuracy and robustness across multiple reasoning benchmarks.
@@ 1 Introduction
LLMs have demonstrated remarkable potential in tasks such as logical reasoning, mathematical proof, and multi-hop question answering. A representative technique to elicit such capabilities is CoT prompting (Wei et al., 2022), which improves interpretability and accuracy by decomposing complex problems into coherent intermediate steps.
Despite these advances, CoT and its extensions fundamentally rely on an autoregressive, step-by-step generation paradigm, where each reasoning step conditions on previously generated outputs Cao et al. (2026). Such sequential dependency increases the susceptibility of the reasoning process to early-stage errors, which may propagate and compound throughout the reasoning trajectory Wang et al. (2025). Moreover, the incremental generation mechanism lacks structured analytical control to regulate redundancy and identify salient reasoning features Chu et al. (2024), resulting in unstable reasoning trajectories and limited interpretability. To alleviate these limitations, a series of structured extensions have been proposed. Approaches such as Tree-of-Thought (ToT) (Yao et al., 2023a), Graph-of-Thought (GoT) (Besta et al., 2024), and Atom-of-Thought (AoT) (Teng et al., 2025) expand the reasoning space beyond single-path CoT and increase reasoning diversity. However, while these methods expand the reasoning space, their ability to mitigate autoregressive dependency remains limited. Interactions across hypotheses are not explicitly coordinated, which may allow errors to accumulate across branches. Other frameworks attempt to enhance efficiency and robustness. ReAct (Yao et al., 2023b) integrates reasoning with acting, AFlow (Zhang et al., 2024) applies Monte Carlo tree search to reasoning workflows, and ReCEval (Prasad et al., 2023) evaluates reasoning chains for correctness and informativeness. While these methods improve task-level performance, they primarily emphasize outcome optimization rather than structural regulation, and thus lack a unified analytical framework for systematically characterizing reasoning chains. As explicit representations of the reasoning process Xia et al. (2025), reasoning chains inherently encode the semantic and logical organization of problem solving. However, without a principled structural perspective to characterize the topology and dependency patterns within reasoning chains, global coordination and consistency regulation remain underdeveloped. As a result, error propagation and redundant reasoning cannot be systematically controlled, limiting the reliability and interpretability of reasoning outcomes.
@@ 5 Conclusion
6 Acknowledgments
This work was partially supported by the National Natural Science Foundation of China under grants 62220106008, and 62572104.